The electric field between two parallel square metal plates is 130 N/C. The plates are 0.85 m on a side and are separated by 3.0 cm. What is the magnitude of the charge on each plate (each plate has equal charge and opposite sign)? Neglect edge effects.


1
Expert's answer
2020-10-21T09:53:44-0400

Assuming that field is perpenicular to the plates, given that the field flow from positive to negative charge, and kwnowing that field is constant we can calculate Q using Gauss's law for this particular case as shown in first equation, and solving for Q.

"EA = \frac{Q}{ε_0}"

"Q = \u03b5_0EA"

"Q = \u03b5_0El^2"

"Q = (8.85 \times 10^{-12})(130)(0.85)^2"

"Q = 8.3 \times 10^{-10} \;C"